OpenStudy (anonymous):

Solve: 4(2x − 1) − 3 = 9(x − 4) −29 32 13 29

OpenStudy (anonymous):

4(2x-1)-3=9(x-4) Step 1. Distribute. => 8x-4-3=9x-36 Step 2. Simplify. => 8x-7=9x-36 Step 3. Combine like terms. => 29=1x Step 4. Divide by the coefficient to get x by itself. Luckily for us, the coefficient is 1 => 29=x